OpenStudy (anonymous):

What is the Area of a circle with a radius of 5. Leave pi in your answer.

OpenStudy (anonymous):

Formula For Area Of A Circle: \[Area=\pi*r^2\] -Substitute 5 in as the radius. \[Area = \pi*5^2\] -Multiply 5 by 5 \[Area=\pi*25\] -Multiply pi (3.14) * 25 \[Area=78.5\] The circle's area is 78.5 units^2

OpenStudy (anonymous):

area =pi x r x r =25pi

OpenStudy (anonymous):

Area of circle=\[\Pi \times r^2\] = \[\pi \times 5^2\] = \[25\]

OpenStudy (anonymous):

25pi